# Multi-scale optimal control for Einstein Telescope active seismic isolation

### v1

@@ Abstract
We present a multi-scale optimal control framework for active seismic isolation in the Einstein Telescope, a third-generation gravitational-wave observatory. Our approach jointly optimizes feedback and blending filters in a cross-coupled opto-mechanical system using a unified cost function based on the “acausal optimum,” which quantifies sensor signal-to-noise ratios across frequencies. This method enables efficient re-optimization under varying sensor configurations and environmental conditions. We apply the framework to two candidate sensing systems: OmniSens—a six-degree-of-freedom inertial isolation system—and BRS-T360, which combines Beam Rotation Sensor (BRS) as an inertial tilt sensor with T360 as a horizontal seismometer. We demonstrate superior low-frequency isolation with OmniSens, reducing platform motion by up to two orders of magnitude near the microseism. The framework allows for ready optimization and projection of sensor noise to metrics relevant to the performance of the instrument, aiding the design of the Einstein Telescope.
@@ 1 Introduction
Current GW detectors are limited by low-frequency noise dominated mostly by a combination of seismic noise and control noise, mainly below 3Hz [6, 2, 5, 1]. The cross couplings—especially tilt-to-translation coupling [13]—play a major role in reducing the sensitivity at low frequency.
@@
Achieving improved low-frequency sensitivity is essential for observing a broader range of astrophysical events over longer timescales, including mergers involving stellar-mass and intermediate-mass black holes. These systems are key to advancing our understanding of fundamental physics and validating diverse theoretical models [12]. The Einstein Telescope (ET) is a proposed third-generation gravitational wave observatory [18, 8], designed to surpass the capabilities of current second-generation detectors—particularly in the low-frequency regime, where seismic and control noise are dominant limiting factors [6, 2, 5, 1]. Addressing this low-frequency noise barrier requires significant upgrades, including enhancements to active isolation systems aimed at suppressing optical component motion at low frequencies [15]. However, designing and modelling a high-performance active isolator is a time-intensive task, especially when accounting for realistic environmental inputs and critical tilt-to-length coupling effects [13]. Control filter design is inherently dependent on the dynamics of sensors, disturbances, and plant. Any modification to these elements necessitates updating control filters to maintain optimal feedback performance and consequently updating noise budgets.
@@
Optimal control has previously been investigated in the design of blending filters for GW seismic isolation platforms [24]. Building on this, an improved weighting function and a generalized plant formulation were later introduced for KAGRA’s active isolation system, incorporating both seismic and sensor noise dynamics [25]. However, in that work, the blending filter is treated independently from the feedback control loop in which it is implemented. In addition, due to the optimization algorithm used, the complexity of the optimal controller produced in these works is in the same order as the plant, making it infeasible to scale to larger plants with cross-couplings and noise dynamics included. A multi-objective optimization approach for the ET suspension system has recently been investigated to identify optimal combinations of actuation filters for the final stage of the suspension, while also quantifying the contribution of Digital to Analog Converter (DAC) noise to the main interferometer axis [23]. This study employs a non-smooth gradient descent algorithm and adopts a more flexible framework that simultaneously considers multiple design criteria, including robustness. It advances optimal control techniques within the context of gravitational wave detectors. Given that control loops often interact, it is crucial to design filters for one loop while accounting for the impact of cross-couplings on the performance of others.
@@
In this paper, we formulate the control of an active isolation platform as a multi-scale optimal-control problem that allows rapid testing and re-optimization for differing input motions and sensor configurations considering cross-couplings and nested loops. At the core of the proposed method is the so-called ”acausal optimum”, denoted by $\xi$, which is formed by taking the minimum across all sources of noise. Its inverse acts as a normalization factor for the platform motion, converting the closed-loop residual motion into a Signal-to-Noise Ratio (SNR). This then enforces the optimizer to use sensors in regions with high SNR and penalizes the loop gain where SNR is low. We construct optimized noise curves for two configurations of sensors: OmniSens [15], a novel 6D inertial isolation system; and a combination of Nanometrics T360 seismometers [16], and Beam Rotation Sensor (BRS) [27] upgraded with interferometric readout [14] here called BRS-T360. This not only sets the stage for further noise propagation and suspension design studies for ET but also provides a novel approach for tackling multi-scale optimization problems abundant within the complex systems of a gravitational wave observatory.
@@ 2 Active isolation configurations
Since for any inertial sensor measuring horizontal motion tilt couples to translation at low frequency [13] we need to have both rotation and translation sensing. In this paper, we explore two configurations: OmniSens and BRS-T360.
@@
The tilt noise performance for the upgraded BRS is constructed from the readout noise and thermal noise of the flexure which gives its proof mass low-frequency tilt compliance. Only these noise sources are considered to provide a fair comparison with the OminSens configuration, despite other limitations to its performance existing in realised implementations to date [27, 26, 21, 20]. Upgrading the readout of the BRS to interferometric sensors [14] is motivated by the successful inclusion of the same in its more compact sibling sensor the Cylindrical Rotation Sensor [19] - here the BRS benefits by a factor of few dues to its larger size reducing the readout noise. Consistent with the other sensors, in both configurations, a transfer function fit to the total noise, the quadrature sum of its readout and flexure thermal noise, of the interferometric readout BRS is used to derive the equivalent sensor noise.
@@ 3 Optimal control
Figure 2: Control schemes for rotation (left, $\theta$) and horizontal translation (right, $x$). White noise inputs $w$ are shaped by noise coloring filters $\eta$ to model real noise sources. The transfer function $T_{\theta\to x}$ captures coupling from rotation to translation. Weighting functions $W$ normalize the closed-loop residual motions $\theta$ and $x$. $G_{\theta}$ and $G_{x}$ are the plant transfer functions, while $K_{\theta}$ and $K_{x}$ are the feedback controllers for rotation and translation, respectively. $H_{\theta}$ and $H_{x}$ are the blending filters, with complementary filters $\bar{H}_{\theta}$ and $\bar{H}_{x}$.
@@
(a) SNR plot for all sources of noise x
(b) SNR plot for all sources of noise $\theta$
Figure 4: OmniSens acausal optimum $\xi$ decomposition for both $\theta$ and x loop. The curves $\eta_{dist}$, $\eta_{disp}$, and $\eta_{sens}$ captures ground disturbance, displacement sensor noise, and inertial sensing noise, respectively.
(a) x loop
(b) $\theta$ loop
@@
To formulate the control design as an optimization problem, we follow the procedure shown in Figure 3. Noise inputs are modeled using noise-coloring transfer functions, which define the acausal optimum $\xi$, formally introduced in Equation 9. These optima are illustrated for both configurations in Figures 4 and 5. The weighting functions are constructed as the inverse of these acausal optimums. Together with the plant transfer functions, these form the generalized plant used in the optimization algorithm. Each step is detailed in the following sections.
The mechanics are designed to minimize cross-couplings, simplifying and clarifying the control design. We assume the coupling from horizontal translation to rotation is negligible. Noise inputs are modelled as white noise multiplied by noise coloring filters $\eta$. These filters are obtained by fitting transfer functions to each noise source’s spectra, with the Laplace variable, s, is omitted for notational simplicity. For the tilt disturbance, the noise coloring filter is:
@@
where here $g_{\rm\theta}$ is the transfer function fit to ground motion and $\tilde{G}_{\rm\theta}$ is the transfer function from ground to the platform. Similarly, for the $x$ direction:
@@
where $\lambda_{x}$ penalizes the open loop motion, encouraging the optimizer to engage the feedback at low frequencies. This ensures the platform remains ground-locked at low frequencies, e.g., for DC actuation to track Earth tide fluctuations.
@@
Where $l$ is the lever arm length from center of rotation. Since the sensor self noise $n(t)$ is uncorrelated from ground-induced rotation on ISI, $g_{\theta}(t)$, i.e., $\langle n(t),g_{\theta}(t)\rangle=0$, the equivalent noise is the quadrature sum of ground and sensor noise given by $\tilde{n}_{\theta}$. The corresponding noise coloring is:
$\eta_{\rm\theta,disp}=\tilde{n}_{\theta}.$ (4)
Analogously, in the $x$ direction:
$\eta_{x,{\rm disp}}=\tilde{n}_{x}.$ (5)
For $\eta_{x,\mathrm{sens}}$ and $\eta_{\theta,\mathrm{sens}}$, which represent the noise coloring filters for the horizontal and tilt inertial sensors respectively, we use transfer function fits, as with other noise coloring filters. These filters are incorporated into the generalized plant control block. For the matrix formulation of the generalized plant, see Appendix C.
@@ 3.1 Loop analysis
$z_{1}=W_{\rm\theta,p}(S_{\theta}w_{\rm\theta,dist}\eta_{\rm\theta,dist}+S_{\theta}K_{\theta}G_{\theta}w_{\rm\theta,disp}\eta_{\rm\theta,disp}\bar{H}_{\rm\theta}+S_{\theta}K_{\theta}G_{\theta}w_{\rm\theta disp}\eta_{\theta,sens}H_{\rm\theta}).\$ (7)
@@
Note the nested structure of the expression above, which features two blending filters. One of these filters is embedded within the primary feedback loop, effectively blending input noise sources with the open-loop response. This layered architecture is why we refer to it as multi-scale optimization, as the inner and outer loops interact in a simultaneous way. The same analysis would apply to the x loop as well.
@@ 3.2 Weighting function
Figure 6: Generalized plant scheme. The generalized plant $P$ includes all transfer functions for noise coloring, weighting filters, and the physical plant. Blue-bordered blocks represent the filters being optimized: $K_{x}$ and $H_{x}$ (control and blending filters for the $x$ loop), and $K_{\theta}$ and $H_{\theta}$ (for the $\theta$  loop). Closed-loop transfer functions from white noise input $\vec{w}$ to output $\vec{z}$ are evaluated using various norms to assess optimality.
In an ideal non-causal case, we can reduce the noise down to the minimum of all sources of noise, namely acausal optimum. Deviations from this non-causal limit measure the achieved performance with our Linear Time Invariant (LTI) filtering scheme. The acausal optimum for both configurations is shown in Figure 4 and 5. We use this acausal optimum with either $H_{2}$ or $H_{\infty}$ norm to form the weighting function. No weighting results in the lowest RMS platform motion over the full frequency band. However, this makes the controller optimization focus on frequencies with high noise and not exploit the frequency bands where the sensors have their best performance. Thereby, a weighting function is introduced to make the optimization aim at the best performance that is feasible with the used sensors, i.e., improve the relative deviation from the theoretically best performance with the given noise sources. This weighting results in substantially lower closed-loop motion at frequencies with high sensing performance at a negligible RMS penalty.
@@
A transfer function fitted to $W_{p}$ is used as the weighting function. Additionally, to ensure uniform emphasis across all frequency decades in the $H_{2}$ norm (on a log-log plot), we apply an extra frequency-dependent weight of the form $f^{-\alpha}$. Setting $\alpha=0.5$ puts the same emphasis across different frequencies. Choosing $0.51>\alpha>$ 0.5 emphasizes low frequencies, while $0.50<\alpha<$ 0.5 emphasizes high frequencies. To implement the resulting fractional-order transfer function, we use the Oustaloup recursive filter method [17]. The complete generalized plant—incorporating all transfer functions, noise coloring filters, and weighting functions—is shown in Figure 6.

### v2

@@ Abstract
We present a multi-scale optimal control framework for active seismic isolation in the Einstein Telescope, a third-generation gravitational-wave observatory. Our approach jointly optimizes feedback and blending filters in a cross-coupled opto-mechanical system using a unified cost function based on the “acausal optimum,” which quantifies sensor signal-to-noise ratios across frequencies. This method enables efficient re-optimization under varying sensor configurations and environmental conditions. We apply the framework to two candidate sensing systems using their modeled sensitivity: OmniSens—a six-degree-of-freedom inertial isolation system—and BRS-T360, which combines Beam Rotation Sensor (BRS) as an inertial tilt sensor with T360 as a horizontal seismometer. We demonstrate superior low-frequency isolation with OmniSens, reducing platform motion by up to two orders of magnitude near the microseism. The framework allows for ready optimization and projection of sensor noise to metrics relevant to the performance of the instrument, aiding the design of the Einstein Telescope.
@@ 1 Introduction
Current GW detectors are limited by low-frequency noise dominated by a combination of seismic noise and control noise, below 3Hz [1, 2, 3, 4]. The cross couplings—especially tilt-to-translation coupling [5]—play a major role in reducing the sensitivity at low frequency.
@@
Achieving improved low-frequency sensitivity is essential for observing a broader range of astrophysical events over longer timescales, including mergers involving stellar-mass and intermediate-mass black holes. These systems are key to advancing our understanding of fundamental physics and validating diverse theoretical models [6]. The Einstein Telescope (ET) is a proposed third-generation gravitational wave observatory [7, 8]. It is designed to surpass the capabilities of current second-generation detectors-particularly in the low-frequency regime, where seismic and control noise are dominant limiting factors [1, 2, 3, 4]. Addressing this low-frequency noise barrier requires significant upgrades, including enhancements to active isolation systems aimed at suppressing optical component motion at low frequencies [9]. However, designing and modelling a high-performance active isolator is a time-intensive task, especially when accounting for realistic environmental inputs and critical tilt-to-translation coupling effects [5]. Control filter design is inherently dependent on the dynamics of sensors, disturbances, and plant. Any modification to these elements necessitates updating control filters to maintain optimal feedback performance and consequently updating the model of the noise limited sensitivity of the detector.
@@
Optimal control has previously been investigated in the design of blending filters for GW seismic isolation platforms [10]. Building on this, an improved weighting function and a generalized plant formulation were later introduced for KAGRA’s active isolation system, incorporating both seismic and sensor noise dynamics [11]. However, in that work, the blending filter is treated independently from the feedback control loop in which it is implemented. Since the blending filter is nested within the feedback loop, this separation implies that the blending filter may be optimal in isolation, while the overall control loop is not globally optimal. In addition, due to the optimization algorithm used, the complexity of the optimal controller produced in these works is in the same order as the plant, making it infeasible to scale to larger plants with cross-couplings and noise dynamics included. A multi-objective optimization approach for the ET suspension system has recently been investigated to identify optimal combinations of actuation filters for the final stage of the suspension, while also quantifying the contribution of Digital to Analog Converter (DAC) noise to the main interferometer axis [12]. This study employs a non-smooth gradient descent algorithm and adopts a more flexible framework that simultaneously considers multiple design criteria, including robustness. It advances optimal control techniques within the context of gravitational wave detectors. Given that control loops often interact, it is crucial to design filters for one loop while accounting for the impact of cross-couplings on the performance of others.
@@
In this paper, we formulate the control of an active isolation platform as a multi-scale optimal-control problem that allows rapid testing and re-optimization for differing input motions and sensor configurations considering cross-couplings and nested loops. The nested loop structure naturally introduces multiple scales, and by optimizing all levels simultaneously, the problem becomes inherently multi-scale. At the core of the proposed method is the so-called ”acausal optimum”, denoted by $\xi$, which is formed by taking the minimum across all sources of noise. Its inverse acts as a normalization factor for the platform motion, converting the closed-loop residual motion into a Signal-to-Noise Ratio (SNR). This normalization factor then enforces the optimizer to use sensors in regions with high SNR and penalizes the loop gain where SNR is low. We construct optimized noise curves for two configurations of sensors: OmniSens [9], a novel 6D inertial isolation system; and a combination of Nanometrics T360 seismometers [13], and Beam Rotation Sensor (BRS) [14] upgraded with interferometric readout [15] here called BRS-T360. This not only sets the stage for further noise propagation and suspension design studies for ET but also provides a novel approach for tackling multi-scale optimization problems abundant within the complex systems of a gravitational wave observatory.
@@ 2 Active isolation configurations
Since for any inertial sensor measuring horizontal motion tilt couples to translation at low frequency [5] we need to have both rotation and translation sensing.
@@
The tilt noise performance for the upgraded BRS is constructed from the readout noise and thermal noise of the flexure which gives its proof mass low-frequency tilt compliance. Only these noise sources are considered to provide a fair comparison with the OminSens configuration, despite other limitations to its performance existing in realised implementations to date [14, 18, 19, 20]. Upgrading the readout of the BRS to interferometric sensors [15] is motivated by the successful inclusion of the same in its more compact sibling sensor the Cylindrical Rotation Sensor [21].
Symbol Description Symbol Description
$G(s)$ Plant transfer function $\xi(f)$ Acausal optimum
$K(s)$ Control filter $S(s)$ Sensitivity function $S(s)=\frac{1}{1+K(s)G(s)}$
$H(s)$ Blending filter (high-pass) $P(s)$ Generalized plant
$\bar{H}(s)$ Complementary blending filter (low-pass) $W_{p}(s)$ Weighting function
$\eta_{\rm dist}(s)$ Disturbance noise coloring filter $w(s)$ White noise input
$\eta_{\rm disp}(s)$ Displacement sensor noise coloring filter $\eta_{\rm sens}(s)$ Inertial sensor noise coloring filter
Table 1: Overview of the commonly used symbols in the paper. Subscripts indicate the degree of freedom (DoF): $\theta$ for rotational loop and $x$ for horizontal loop. For example, $K_{\theta}$ and $\eta_{\theta,\mathrm{dist}}$ correspond to the $\theta$ loop, while $K_{x}$ and $\eta_{x,\mathrm{dist}}$ correspond to the x loop.
@@ 3 Optimal control
Figure 2: Control schemes for rotation (left, $\theta$) and horizontal translation (right, $x$). White noise inputs $w$ are shaped by noise coloring filters $\eta$ to model real noise sources. The transfer function $T_{\theta\to x}$ captures coupling from rotation to translation. Weighting functions $W$ normalize the closed-loop residual motions $\theta$ and $x$ in order to produce performance outputs $z_{1}$ and $z_{2}$. $G_{\theta}$ and $G_{x}$ are the plant transfer functions, while $K_{\theta}$ and $K_{x}$ are the feedback controllers for rotation and translation, respectively. $H_{\theta}$ and $H_{x}$ are the blending filters, with complementary filters $\bar{H}_{\theta}$ and $\bar{H}_{x}$. $\eta_{\mathrm{dist}}$, $\eta_{\mathrm{disp}}$, and $\eta_{\mathrm{sens}}$ represent the disturbance, displacement-sensor, and inertial-sensor noise coloring filters, respectively; the subscript $\theta$ or $x$ indicates the rotational or horizontal loop.
@@
(a) $\theta$ loop.
(b) x loop.
Figure 4: OmniSens acausal optimum $\xi$ decomposition for both $\theta$ and x loop. The acausal optimum represents the global minimum across all noise curves, as defined in Eq. 7. The curves $\eta_{dist}$, $\eta_{disp}$, and $\eta_{sens}$ captures ground disturbance, displacement sensor noise, and inertial sensing noise, respectively.
(a) $\theta$ loop.
(b) x loop.
@@
This section presents a multi-scale control design that combines both $H_{\infty}$ (angular) and $H_{2}$ (translation) norms within a generalized plant framework. Weighting functions derived from the signal-to-noise ratio (SNR) quantify deviation from the acausal optimum—representing the theoretical performance limit—so that optimization targets realistic, sensor-limited performance rather than uniform RMS minimization. In the $H_{2}$ norm, an additional logarithmic weighting proportional to $f^{-\alpha}$ ($\alpha=0.5$) ensures equal contribution per frequency decade. Optimization is performed using a subgradient descent algorithm rather than Riccati-based synthesis to enable reduced-order, numerically robust controllers. Although subgradient descent sacrifices strict optimality guarantees, it circumvents the numerical conditioning problems that prevented Riccati methods from converging to true optima in practical, multi-loop implementations.
@@
Optimal control assumes all noise inputs are white. Noise coloring filters shape these inputs to reflect realistic disturbances and sensor noise, while weighting functions emphasize specific frequency ranges in the plant output to guide the optimization toward desired performance. The combination of plant, noise filters, and weighting functions forms the generalized plant, which provides the basis for controller optimization. For a detailed introduction to optimal control, see [22].
To formulate the control design as an optimization problem, we follow the procedure shown in Figure 3. Noise inputs are modeled using noise-coloring transfer functions, which define the acausal optimum $\xi$, formally introduced in Equation 7. These optima are illustrated for both configurations in Figures 4 and 5. Overall, OmniSens demonstrates superior performance owing to its lower thermal noise and increased sensitivity to rotation at low frequencies. In addition, the interferometric readout of the OmniSens reference mass for translation provides improved sensitivity compared to that of the T360.
The weighting functions are constructed as the inverse of these acausal optimums. Together with the plant transfer functions, these form the generalized plant used in the optimization algorithm. Each step is detailed in the following sections.
The mechanics are designed to minimize cross-couplings, simplifying and clarifying the control design. We assume the coupling from horizontal translation to rotation is negligible. Noise inputs are modelled as white noise multiplied by noise coloring filters $\eta$. These filters are obtained by fitting transfer functions to each noise source’s spectra. The Laplace variable, s, is omitted for notational simplicity. For the tilt disturbance, the noise coloring filter is:
@@
where here $g_{\rm\theta}$ is the transfer function fit to ground motion spectra and $\tilde{G}_{\rm\theta}$ is the transfer function from ground to the platform. Similarly, for the $x$ direction:
@@
where $\lambda_{x}$ is a low-pass filter with a pole set below table resonance which penalizes the open loop motion, encouraging engagement of feedback at low frequencies. This ensures that the controller maintains high loop gain at low frequencies and provides a summing junction for external inputs, such as observatory drift control signals.
@@
Where $l$ is the lever arm length from center of rotation and $\theta(t)$ is table tilt. Since the sensor self noise $n(t)$ is uncorrelated from ground rotation, $g_{\theta}(t)$, i.e., $\langle n(t),g_{\theta}(t)\rangle=0$, the equivalent noise is the quadrature sum of ground and sensor noise which is used in the noise coloring filter, $\eta_{\rm\theta,disp}$. The same procedure applies in the x direction, yielding $\eta_{x,{\rm disp}}$.
For $\eta_{x,\mathrm{sens}}$ and $\eta_{\theta,\mathrm{sens}}$, which represent the noise coloring filters for the horizontal and tilt inertial sensors respectively, we use transfer function fits, as with other noise coloring filters.
For implementation, we take the control blocks shown in 2 and rearrange the diagram such that all the white-noise inputs $w$ appear on the left and all performance outputs $z$ on the right. We replace complementary filters in the scheme by applying $\bar{H}=1-H$. This allows us to only optimize for the $H$ and derive the other by using the complementarity property [11]. We isolate four filters to be optimized ($K_{x}$, $H_{x}$, $K_{\theta}$, and $H_{\theta}$) and everything which is left is part of the generalized plant shown in 6. We can express the obtained input-output system using a matrix. This matrix can be used in Matlab to define the generalized plant. For the matrix formulation of the generalized plant, see Appendix C.
@@ 3.1 Loop analysis
$z_{1}=W_{\rm\theta,p}(S_{\theta}w_{\rm\theta,dist}\eta_{\rm\theta,dist}+S_{\theta}K_{\theta}G_{\theta}w_{\rm\theta,disp}\eta_{\rm\theta,disp}\bar{H}_{\rm\theta}+S_{\theta}K_{\theta}G_{\theta}w_{\rm\theta disp}\eta_{\theta,sens}H_{\rm\theta}).\\$ (5)
@@
Note the nested structure of the expression above, which features two blending filters. One of these filters is embedded within the primary feedback loop, effectively blending input noise sources with the open-loop response. This layered architecture is why we refer to it as multi-scale optimization, as the inner and outer loops interact in a simultaneous way. The same analysis applies to the x loop as well.
@@ 3.2 Weighting function
Figure 6: Generalized plant scheme. The generalized plant $P$ includes all transfer functions for noise coloring, weighting filters, and the physical plant. $K_{x}$ and $H_{x}$ (control and blending filters for the $x$ loop), and $K_{\theta}$ and $H_{\theta}$ (for the $\theta$ loop) are being optimized. Closed-loop transfer functions from white noise input $\vec{\omega}=[w_{\theta,\rm dist}\ w_{\theta,\rm disp}\ w_{\theta,\rm sens}\ w_{x,\rm dist}\ w_{x,\rm disp}\ w_{x,\rm sens}]^{T}$ to output $\vec{z}=[z_{1}\ z_{2}]^{T}$ are evaluated using various norms to assess optimality.
In an ideal non-causal case, we can reduce the noise down to the minimum of all sources of noise, namely the acausal optimum. Deviations from this non-causal limit measure the achieved performance with our Linear Time Invariant (LTI) filtering scheme. The acausal optimum for both configurations is shown in Figure 4 and 5. We use this acausal optimum with either $H_{2}$ or $H_{\infty}$ norm to form the weighting function [22]. No weighting results in the lowest RMS platform motion over the full frequency band. However, this makes the controller optimization focus on frequencies with high noise and not exploit the frequency bands where the sensors have their best performance. Thereby, a weighting function is introduced to make the optimization aim at the best performance that is feasible with the used sensors, i.e., improve the relative deviation from the theoretically best performance with the given noise sources. This weighting results in substantially lower closed-loop motion at frequencies with high sensing performance at a negligible RMS penalty.
@@
The weighting function in the Laplace domain, $W_{p}(s)$, is obtained by fitting a transfer function to the frequency-domain representation, $W_{p}(f)$. Additionally, to ensure uniform emphasis across all frequency decades in the $H_{2}$ norm (on a log-log plot), we apply an extra frequency-dependent weight of the form $f^{-\alpha}$. Setting $\alpha=0.5$ (the value used in this study) puts the same emphasis across different frequencies. Choosing $1>\alpha>0.5$ emphasizes low frequencies, while $0<\alpha<0.5$ emphasizes high frequencies. To implement the resulting fractional-order transfer function, we use the Oustaloup recursive filter method [23]. The complete generalized plant—incorporating all transfer functions, noise coloring filters, and weighting functions—is shown in Figure 6.